Question 57 1 pts Herring Corporation has operating income of $235,000 and a 40% tax rate. The firm has short-term debt of $115,000, long-term debt of $321,000, and common equity of $436,000. What is its return on invested capital? 14.92% O 15.50% 16.17% O 17.42% 18.27% s Previous Next F3 F4 FS 1 F6 F7 F8 19 5
Answer #1

Answer : "16.17%"

=> Invested Capital = STD + LTD + EQUITY = 115000+321000+436000 = 872000

Operating Income (IBT) = 235000

Net Income = OI * tax rate = 235000 * (1 - 0.4) = $141000

Return on invested Capital = 141000/872000 = 16.17%